OpenStudy (anonymous):

Arrange these numbers from least to greatest: -5/6, -4/9, -3/5, -4/7 Does anybody have any idea how to do this? the negative signs are making me confused

OpenStudy (anonymous):

-5/6 = -0.83... -4/9 = -0.44... -3/5 = -0.6 -4/7 = -0.57

OpenStudy (anonymous):

So would it be -3/5, -4/9, -4/7. -5/6?

OpenStudy (anonymous):

Nope, whats the least number here?

OpenStudy (anonymous):

-83?

OpenStudy (anonymous):

No...your answer is: -4/9, -4/7, -3/5, -5/6.

OpenStudy (anonymous):

Sorry I'm not the best at math I struggle little :/

OpenStudy (anonymous):

It's alright :)

OpenStudy (anonymous):

Thank you for the help :)

OpenStudy (anonymous):

I always say think of a number line if you get stuck doing these arithmetic problems.

OpenStudy (anonymous):

wait wait

OpenStudy (anonymous):

I put it from greatest to least, just put it the other way for least to greatest :P

OpenStudy (anonymous):

I said it backwards XD. Sorry for being dumb >.<

OpenStudy (anonymous):

ohh ok Thank you that's a good idea, I'll try that next time :) And it's okay your not dumb it was just a mistake, I make mistakes ALL the time XD

OpenStudy (anonymous):

Haha, I always thinking one thing and write another >_>.

OpenStudy (anonymous):

am*

OpenStudy (anonymous):

Same here :)

OpenStudy (anonymous):

It's probably due to the lack of sleep, haha sorry for going on like that, but yes make a number line it will help you a lot!

OpenStudy (anonymous):

okay thx :D

OpenStudy (anonymous):